OpenStudy (anonymous):

ASKING FOR ANYONE'S HELP! WILL GIVE MEDAL AND I WILL FAN YOU!!!! have been stuck since 7:00 p.m. Question 23 Given the geometric sequence where a1 = 1 and the common ratio is 6, what is the domain for n? All integers where n ≥ 1 All integers All integers where n ≥ −1 All integers where n ≥ 0 Question 24 Find the sum of a finite geometric sequence from n = 1 to n = 5, using the expression −3(4)n − 1. -1,023 1,223 1,023 −4,374

OpenStudy (anonymous):

Question 25 What is the 7th term of the geometric sequence 4, −20, 100, …? −312,500 −12,500 62,500 1,562,500 Question 26 What is the sum of the arithmetic sequence 153, 139, 125, …, if there are 22 terms? 132 154 176 198 Question 27 What is the sum of the geometric sequence −1, 6, −36, … if there are 7 terms? −39,991 −6,665 6,665 39,991 Question 28 Braxton loves to download new music. He originally had 41 songs but plans to purchase 4 new songs each week. He wants to know how many songs he will have after 110 weeks. Which equation should he use? a110 = 41 + 4(110 − 1) a110 = 41 + 4(110 + 1) a110 = 110 + 4(41 − 1) a110 = 110 + 4(41 + 1) Question 29 What is the sum of the geometric sequence 3, 15, 75, … if there are 7 terms? 39,062 58,593 195,312 292,968 Question 30 Find the 12th partial sum of the summation of negative 7i plus 22, from i equals 1 to infinity. 25 40 -348 72 Question 31 Given the geometric sequence an = 3(−4)n − 1, which of the following values for n lies in the appropriate domain for n? n = 0 n = −1 n = 1 They all lie in the domain Question 32 The surface area of an oil spill gets 131% larger every day, represented by the function s(x) = (1.31)x − 1. On the first day, it covered an area of 21 square meters. Which function would be used to find the area of the oil spill on the 47th day? s(x) = (1.31)21 − 1 s(x) = 21(1.31)x − 1 s(x) = (1.31)x − 21 s(x) = 1.31(21)x − 1

OpenStudy (anonymous):

pick one

OpenStudy (anonymous):

Question 23 Given the geometric sequence where a1 = 1 and the common ratio is 6, what is the domain for n? all integers \(n\geq 1\) for that one

OpenStudy (anonymous):

Thank you hun. (:

OpenStudy (anonymous):

ready for the next one?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sum+n+%3D+1+to+5+%28-3*4^ {n-1}%29

OpenStudy (anonymous):

Question 25 What is the 7th term of the geometric sequence 4, −20, 100, formula must look like \(4\times (-5)^{n-1}\) so 7th term is \[4\times (-5)^6\]

OpenStudy (anonymous):

uhm... im not sure but i dont think something came up right for 24....

OpenStudy (anonymous):

look at the link i sent it has the answer there for number 25 the answer is \[4(-5)^6=62500\] http://www.wolframalpha.com/input/?i=4%28-5%29^5

OpenStudy (anonymous):

#24 the answer was \(-1023\)

OpenStudy (anonymous):

how we doing, okay?

OpenStudy (anonymous):

yeah... (:

OpenStudy (anonymous):

where we at #26?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

can you do me a favor and repost just 26, i can't keep scrolling up so far i get lost

OpenStudy (anonymous):

Question 26 What is the sum of the arithmetic sequence 153, 139, 125, …, if there are 22 terms? 132 154 176 198

OpenStudy (anonymous):

ok here \(n=22\) \(a_1=153, a_n=21\times (-14)+153=-141\) we need \[\frac{n}{2}(a_1+a_n)=\frac{22}{2}(153-141)=11\times 12=132\]

OpenStudy (anonymous):

Question 27 What is the sum of the geometric sequence −1, 6, −36, … if there are 7 terms?

OpenStudy (anonymous):

\[-1+6+6(-6)+6(-6)^2+6(-6)^3+6(-6)^4+6(-6)^5+6(-6)^6=-39991\] http://www.wolframalpha.com/input/?i=-1%2B6%2B6%28-6%29%2B6%28-6%29^2%2B6%28-6%29^3%2B6%28-6%29^4%2B6%28-6%29^5

OpenStudy (anonymous):

DO YOU KNOW HOW FREAKING AWESOME YOU ARE!?!?!

OpenStudy (anonymous):

Question 28 Braxton loves to download new music. He originally had 41 songs but plans to purchase 4 new songs each week. He wants to know how many songs he will have after 110 weeks. Which equation should he use?

OpenStudy (anonymous):

\[a_{110}=41+4(110-1)\]

OpenStudy (anonymous):

lol thanks!

OpenStudy (anonymous):

What is the sum of the geometric sequence 3, 15, 75, … if there are 7 terms

OpenStudy (anonymous):

youre welcome, but m so serious!! If it werent for you id still be stuck on #23 no joke!!

OpenStudy (anonymous):

\[3+3\times 5+3\times 5^2+3\times 5^3+3\times 5^4+3\times 5^5+3\times 5^6\] this is a formula for this, but with only 7 numbers you can just grind it til you find it, use a calculator

OpenStudy (anonymous):

alright. (:

OpenStudy (anonymous):

by calculator i mean wolfram http://www.wolframalpha.com/input/?i=3%2B3 \times+5%2B3\times+5^2%2B3\times+5^3%2B3\times+5^4%2B3\times+5^5%2B3\times+5^6

OpenStudy (anonymous):

looks like the answer is \(58593\) is that a choice? would you like to know the formula?

OpenStudy (anonymous):

it is indeed a choice! (: and sure!

OpenStudy (anonymous):

\[a\left (\frac{1-r^n}{1-r}\right)\] where \(a\) is the first term and \(r\) is the common ratio

OpenStudy (anonymous):

ohhh mkkayy

OpenStudy (anonymous):

so for example for \(3, 15, 75,\) the first term is \(3\) and the common ratio is \(5\) you want the sum of the first \(7\) terms, compute \[3\left(\frac{1-5^7}{1-5}\right)\] or \[3\left(\frac{5^7-1}{5-1}\right)\]

OpenStudy (anonymous):

look, it is the same number! http://www.wolframalpha.com/input/?i=3%285^7-1%29%2F4

OpenStudy (anonymous):

where we at?

OpenStudy (anonymous):

Question 30 Find the 12th partial sum of the summation of negative 7i plus 22, from i equals 1 to infinity yes?

OpenStudy (anonymous):

uhhh i think 29? Question 29 What is the sum of the geometric sequence 3, 15, 75, … if there are 7 terms? 39,062 58,593 195,312 292,968

OpenStudy (anonymous):

we just did that one, twice it is \(58,593\) for that one lets move on

OpenStudy (anonymous):

my bad.. im so lost haha so i guess 30

OpenStudy (anonymous):

Question 30 Find the 12th partial sum of the summation of negative 7i plus 22, from i equals 1 to infinity this question was written by an idiot what they want is \[\sum_{i=1}^{12}-7i+22\]

OpenStudy (anonymous):

there is a formula for this too, but the answer is \(-282\) unlose yourself, you don't want to put the right answer in the wrong spot!!

OpenStudy (anonymous):

haha yeah.

OpenStudy (anonymous):

and thats not an answer choice....

OpenStudy (anonymous):

unfortunately that does not seem to be an answer choice let me check it again

OpenStudy (anonymous):

it does say this Find the 12th partial sum of the summation of negative 7i plus 22 right? or maybe \(-7i+22\) right?

OpenStudy (anonymous):

check it because i am going to stick with my answer, unless there is a typo there

OpenStudy (anonymous):

25 40 -348 72 these are answer choices....

OpenStudy (anonymous):

or just pick the one negative number, because it is definitely negative, even though that answer is wrong

OpenStudy (anonymous):

haha okay! (:

OpenStudy (anonymous):

now we are at 31?

OpenStudy (anonymous):

yes only 2 to go!! (:

OpenStudy (anonymous):

Given the geometric sequence an = 3(−4)n − 1, which of the following values for n lies in the appropriate domain for n? n = 0 n = −1 n = 1 this is an easy one only \(n=1\) is in the domain

OpenStudy (anonymous):

YOUR ARE FOREAL FOREAL AMAZINGLY WONDERFUL!!!

OpenStudy (anonymous):

Question 32 The surface area of an oil spill gets 131% larger every day, represented by the function \( s(x) = (1.31)^{x − 1}\). On the first day, it covered an area of 21 square meters. Which function would be used to find the area of the oil spill on the 47th day?

OpenStudy (anonymous):

thanks (blush)

OpenStudy (anonymous):

Haha no, Thank YOU!

OpenStudy (anonymous):

again this question was written by an idiot, you need a number out front want they want i think is \[s(x)=21\times(1.31)^{x-1}\] there are so many things wrong with this question i can't begin to explain them all, but that is the answer they are looking for who writes this jibberish?

OpenStudy (anonymous):

haha thank you! (:

OpenStudy (anonymous):

you done now?

OpenStudy (anonymous):

i just submitted lets see how we did!! (:

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

probably missed the one with no right answer...

OpenStudy (anonymous):

we missed this one.... Question 27 (Worth 1 points) (08.04) What is the sum of the geometric sequence −1, 6, −36, … if there are 7 terms? −39,991 −6,665 6,665 39,991 Points earned on this question: 0

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

we actually got the one that had no right answer correct... (:

OpenStudy (anonymous):

ho ho ho well missing one is ok i guess

OpenStudy (anonymous):

now it is friday night, i am old and going to be you should go out and do something more fun than math!! (for sure) good luck

OpenStudy (anonymous):

yes!!!! waaayyyyyyy better than my trying to answer would have been hahaha!!! thank you soo much!!!!! YOURE AMAZINGGG!!!!

OpenStudy (anonymous):

"old and going to bed" is what i meant you are quite welcome

OpenStudy (anonymous):

how old are you? lol

OpenStudy (anonymous):

older than black pepper who was born dec 15?

OpenStudy (anonymous):

WHAT?^^^

OpenStudy (anonymous):

i figured "ilovehim121511" meant you loved someone born dec 15 2011 but i could be wrong, it was just a guess

OpenStudy (anonymous):

oh lol no, thats mine and my boyfriends date. (: but no foreal.. how old are u?

OpenStudy (anonymous):

older than you and your boyfriend put together older than auld lang syne so old that when i went to school they didn't teach history ....

OpenStudy (anonymous):

gnight, ttyl

OpenStudy (anonymous):

omg lol nooo foreall!!!!!

OpenStudy (elizabeths):

haha, well that means he doesn't wants to share his age.

OpenStudy (anonymous):

yeah.. i think so. .-.

OpenStudy (elizabeths):

Well, they are some people who don't like sharing, Like me (;

OpenStudy (anonymous):

yeah. dont matter to me,hah

OpenStudy (anonymous):

but u look like u could be anywhere from 14-17 soo yeah

OpenStudy (elizabeths):

14? Nope! nor 17 (;

OpenStudy (anonymous):

ima go with 15.

OpenStudy (elizabeths):

No. Older than 17 =P

OpenStudy (anonymous):

mmh, im sticking with 15...

OpenStudy (elizabeths):

lol I'm not 15 hun, I don't even look like 15.

OpenStudy (anonymous):

haha sweet heart you do (; kay byyeee now!

OpenStudy (elizabeths):

bye and thanks (;